A food shop mixes nuts selling at $16/lb with fruit selling at $3/lb. They want a mixture of 7 lb. to sell at $12/lb. How many pounds of each should they use?
I tried: $16(n)+ $3(7 - n) = $12(7) and got n = 4 and 9/13 lb. so
f = 2 and 4/13 lb. Is this right?

Answer by ptaylor(2198)

You and I arrived at the same equation to solve but I think that you made a mistake in solving the equation.
Let x=Number of lbs of nuts
Then 7-x=Number of lbs of fruit
Now we know that the amount of $ collected from selling the nuts(x)($16) and fruit(7-x)($3) separately would have to equal the amount of $ collected from selling the final mixture (7)($12). Thus, our equation to solve is:
$16x+(7-x)($3)=7($12) Simplifying, we get:
16x+21-3x=84 or
13x=63
x=4 and 11/13 lbs of nuts
7-x=7-4 and 11/13 = 2 and 2/13 lbs of fruit
Hope this helps-----ptaylor